OpenStudy (anonymous):

convert 4x-y=-1 to the formula y=mx+b also convert -x+y=x-5

OpenStudy (anonymous):

y = 4x+1 y = 2x-5

OpenStudy (anonymous):

but there is a way of doing it

OpenStudy (anonymous):

you just isolate y on one side

OpenStudy (anonymous):

ok o hope that is it

OpenStudy (anonymous):

ok what about -x+y=x-5

OpenStudy (anonymous):

yeah you just move the x to the other side so it becomes y = x + x -5 which is y = 2x-5

OpenStudy (anonymous):

oh well that make sense